OpenStudy (anonymous):

i need help with questions 5,6,9,10,11

OpenStudy (anonymous):

@KamiBug

OpenStudy (anonymous):

Can you please help me?

OpenStudy (anonymous):

@C.A.M

OpenStudy (anonymous):

i only know nimber 9 its negative sorry i couldnt help more

OpenStudy (anonymous):

@YanaSidlinskiy

OpenStudy (anonymous):

@mathmath333

OpenStudy (mathmath333):

for \(5.\) see the points https://www.desmos.com/calculator/vb8imkbqow the points seems to be in straight line so choose any two ponits and form the equation of line that will be the equation of line of best fit

OpenStudy (mathmath333):

from the equation of line of best fit in the form \(y=mx+b\), \(m\) will be the slope we see that that as x(years) increase the value of y(dollars) decrease so , slope represents the decrease in value of the car over the years

OpenStudy (anonymous):

so what would i put for the slope

OpenStudy (anonymous):

@johnweldon1993

OpenStudy (mathmath333):

consider two points \(\large \begin{align} \color{black}{ (x_1,y_1)=(0,23)\hspace{.33em}\\~\\ (x_2,y_2)=(1,20)\hspace{.33em}\\~\\ y-y_1=m(x-x_1)\hspace{.33em}\\~\\ m=\dfrac{x_2-x_1}{y_2-y_1}\hspace{.33em}\\~\\ }\end{align}\)

OpenStudy (mathmath333):

where \(\large\tt \begin{align} \color{black}{ \normalsize \text{m is slope }\hspace{.33em}\\~\\ }\end{align}\)